5. Your friend Kevin is evaluating the stocks of North Great Timber Company. North Great Timber Company expects to have an EPS of $6/share next year and pay a dividend of $1.50 a share. You expect that the firms’ ROE (meaning the return on both old and new investments) will stay at 5% in the future and its payout ratio will remain unchanged. According to Kevin’s analysis of the firms’ past stock return, the firm’s beta is 1.25. He also expects that the market return will be 6% and the T-bill rate will be 2%. Currently the market price of the stock is $50/share. (10pts)

  1. (1) What is the expected growth rate of the firms’ dividend?

  2. (2) According to CAPM, what is the appropriate required rate of return on the stock?

  3. (3) Please estimate the fair value of the stock using DDM. Use the required return derived

    from CAPM as the discount rate in the DDM model.

  4. (4) Can you make an investment recommendation to your friend?
Answer #1

(1) What is the expected growth rate of the firms’ dividend?
=RoE*(1-payout ratio)=5%*(1-1.5/6)=3.75%

(2) According to CAPM, what is the appropriate required rate of return on the stock?
=risk free rate+beta*(market return-risk free)=2%+1.25*(6%-2%)=7%

(3) Please estimate the fair value of the stock using DDM. Use the required return derived from CAPM as the discount rate in the DDM model.
=Expected Dividend/(required return-growth rate)=1.5/(7%-3.75%)=46.15384615

(4) Can you make an investment recommendation to your friend?
Sell or dont buy the share as the stock is overvalued because market price is more than the fair value